# Learning from Negative Samples in Generative Biomedical Entity Linking

### v1

@@ 5 Analysis
We conduct comprehensive analyses to evaluate the effectiveness of our method.
@@ 5.1 Ablation Study
5.1.1 Effect of Pre-training
@@
Table 2 highlights the effectiveness of ANGEL’s pre-training by comparing other pre-training methods. BART, pre-trained using a standard language modeling objective but not specifically tailored for BioEL tasks, shows the lowest performance. In contrast, GenBioEL, pre-trained using synonyms from UMLS in a similar manner to our positive-only training, achieves commendable performance despite not being trained on the target data. When ANGEL’s negative-aware training is incorporated into GenBioEL, its performance significantly improves, with gains of 16.6% on BC5CDR and 10.9% on AAP. These performance improvements also carry over to the fine-tuning stage on the target datasets. Our method still leads to an increase in performance, with a 0.2% improvement on both BC5CDR and AAP compared to GenBioEL.
| BC5CDR | AAP | |||
| Pre-trained Model | PT | FT | PT | FT |
| BART | 0.8 | 93.0 | 15.6 | 88.7 |
| GenBioEL | 33.1 | 93.1 | 50.6 | 89.3 |
| + ANGEL (Ours) | 49.7 | 93.3 | 61.5 | 89.5 |
5.1.2 The Number of Synonyms
To evaluate the impact of incorporating multiple synonyms during fine-tuning, we conducted experiments by varying the number of synonyms associated with each mention, testing with 1, 3, and 5 synonyms. As shown in Figure 3, the average performance improves when using three synonyms compared to just one, but it declines when expanding to five synonyms. When the number of available synonyms is less than the specified number (e.g., fewer than 3 or 5 synonyms), all available synonyms are used to ensure maximum diversity in learning. Therefore, unlike GenBioEL, which utilized only the top-1 synonym, our approach incorporated up to the top-3 synonyms per mention, which proved to be optimal.
| Method | BC5CDR | AAP |
| ANGEL (Ours) | 94.5 | 90.2 |
| Prediction-based | ||
| TF-IDF-based | 94.4 (-0.1) | 90.0 (-0.2) |
| Pairs | ||
| All Possible Pairs | 94.0 (-0.5) | 90.0 (-0.2) |
| within Top-5 | ||
| Top-10 Predictions | 94.4 (-0.1) | 89.9 (-0.3) |
5.1.3 Selection of Positive and Negative Pairs
Table 3 presents various methods for constructing positive-negative pairs in negative-aware training. We investigated the effects of three different aspects: negative sampling techniques, the ranking of negative samples, and top-k selection, using the GenBioEL model on the BC5CDR and AAP datasets. First, when we modified our approach to extract negative samples based on TF-IDF, similar to the method used during pre-training, performance declined by 0.1% on the BC5CDR and by 0.2% on the AAP. This indicates that our approach, where the model learns from its errors, is more effective than relying on TF-IDF similarity. Additionally, when we substituted our negative selection method—where negative samples are ranked higher than positive ones—with an approach that includes all positive-negative pairs regardless of rank, the performance dropped by 0.5% on the BC5CDR and by 0.2% on the AAP. Finally, increasing the top-k selection from the top-5 to the top-10 predictions resulted in a performance decline of 0.1% on the BC5CDR and 0.3% on the AAP. This suggests that rather than simply increasing the top-k, it may be more effective to focus on utilizing the positive and negative samples from top-ranked results.
@@ 5.2 Case Study
Figure 4 illustrates the predictions of SapBERT, GenBioEL, and ANGEL. In the first example, the mention ‘ASPD’ is an abbreviation for ‘antisocial personality disorder’ (also known as ‘dissocial personality disorder’). SapBERT incorrectly predicts ‘ASP’ due to the similarity in surface form. GenBioEL struggles to distinguish between correct entity names and those containing the words ‘personality disorder’. In contrast, our model successfully identifies the correct entities, without being misled by false entity names that contain overlapping terms. The second example involves the mention ‘basaglar,’ a biosimilar medication that contains insulin glargine, a long-acting insulin. The challenge here arises from the fact that product names can differ significantly from the biomedical terms used to describe their active ingredients. This discrepancy leads to failures in both SapBERT and GenBioEL, as they struggle to connect the brand name to its corresponding biomedical entity. Nevertheless, our model successfully identifies the correct entity, showcasing its ability to handle such complex cases effectively. In the final example, our method was less effective. For the mention of ‘glutamine,’ neither SapBERT nor GenBioEL identified the correct answer, but they did rank ‘Glutamic acid,’ the correct entity, within the top 5 candidates. Our model, however, ranked the correct answer slightly lower. Consequently, while our model shows a notable improvement in top-1 accuracy, the increase in top-5 accuracy is relatively modest in some datasets. The effectiveness of our method also varies across different datasets. We discuss this limitation in more detail in the Appendix, noting that such cases are an area for further exploration.
@@ 6 Conclusions
In this study, we discussed the importance of negative samples in training generative BioEL models and introduced ANGEL, the first framework in this field to effectively incorporate negative-aware training into a generative model. Our models demonstrated the ability to learn subtle distinctions between entities with similar surface forms and contexts. Experimental results showed that ANGEL outperformed existing similarity-based and generative models, achieving the best performance across five public BioEL datasets.
Versatility is one of the key strengths of our approach, as it can be effectively applied to both the pre-training and fine-tuning stages and is model-agnostic, making it suitable for any type of generative model. Our experiments have consistently demonstrated the significant benefits of incorporating our method in both pre-training and fine-tuning phases, with notable performance improvements of 0.9%, 1.4%, and 1.7% for BART, BioBART, and GenBioEL, respectively In future research, applying our method to decoder-only models Luo et al. (2022) with larger parameter sets Touvron et al. (2023); Chen et al. (2023), in addition to encoder-decoder models, could be particularly valuable in domains where precision is paramount.
@@ Acknowledgements
This research was supported by (1) the National Research Foundation of Korea (NRF2023R1A2C3004176, RS-2023-00262002), (2) the Ministry of Health & Welfare, Republic of Korea (HR20C0021(3)), (3) ICT Creative Consilience Program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-2024-20200-01819), (4) Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2022-00155911, Artificial Intelligence Convergence Innovation Human Resources Development (Kyung Hee University)) and (5) the MSIT (Ministry of Science and ICT), Korea, under the ITRC (Information Technology Research Center) support program (IITP-2024-RS-2024-00438239) supervised by the IITP (Institute for Information & Communications Technology Planning & Evaluation).
Appendix A Details of Pre-training
Our pre-training process follows the KB-Guided Pre-training strategy outlined in the GenBioEL framework Yuan et al. (2022b). We define clause templates to generate synthetic training examples by incorporating synonyms and definitions from the KB. Specifically, we select pairs of synonyms and from the set , along with a definition corresponding to the identifier . The synonyms and definitions are integrated into one of the two predefined clause templates as follows:
| or | ||
The input for the decoder is “” and the output should be “”. When no definitions are available in the KB, we construct using alternative synonyms. For concepts with only two synonyms, and are used as the synonyms, with being the same as . For concepts with only one synonym, , , and are the same, resulting in a straightforward sentence. This strategy effectively creates simulated contexts for the model to learn from, improving its ability to generalize to new entities not included in the downstream datasets. The complete list of templates used in this process is provided in Table A.
| Template | |
| Encoder Side | Decoder Side |
| < is defined as > | is |
| < is described as > | |
| < are the definitions of > | |
| < describe > | |
| < define > | |
| < are the synonyms of > | |
| < indicate the same concept as > | |
| < has synonyms such as > | |
| < refers to the same concepts as > | |
| < is > | |
| < is the same as > | |
| < is > | |
| < is the same as > | |
Appendix B Datasets
@@
| Entity names | 108,071 | 809,929 | 904,798 | 3,381 | 203,282 |
| Identifiers | 14,967 | 268,162 | 350,830 | 1,036 | 25,419 |
@@
Table B presents the statistics of the five datasets used, along with their corresponding target knowledge bases.
NCBI-disease Doğan et al. (2014)
@@
The NCBI-disease dataset contains 793 PubMed abstracts annotated with 6,892 disease mentions that are mapped to 790 unique disease concepts using the MEDIC ontology Davis et al. (2012). MEDIC is a medical dictionary that integrates disease concepts, synonyms, and definitions from both MeSH Lipscomb (2000) and OMIM Hamosh et al. (2004), encompassing a total of 9,700 unique disease entities. This dataset is primarily used for disease recognition and concept normalization tasks.
BC5CDR Li et al. (2016b)
@@
The BC5CDR dataset includes 1,500 PubMed abstracts with 4,409 chemical entities, 5,818 disease entities, and 3,116 chemical-disease interactions. All annotated entities are mapped to the MeSH ontology Lipscomb (2000), which is a subset of UMLS Bodenreider (2004). This dataset is widely used for biomedical entity recognition and interaction studies. To fit the purpose of our study, we use only the chemical and disease annotations and discard the interaction annotations.
COMETA Basaldella et al. (2020)
@@
COMETA focuses on layman medical terminology, compiled from four years of content across 68 health-related subreddits. This dataset consists of 20,000 biomedical entity mentions annotated with concepts from SNOMED CT Chang and Mostafa (2021). It is utilized for the normalization of consumer health expressions into standardized medical terminologies.
AskAPatient (AAP) Limsopatham and Collier (2016b)
@@
The AskAPatient dataset contains 8,662 phrases from social media language, each mapped to medical concepts from SNOMED CT Chang and Mostafa (2021). This dataset does not include contextual information, meaning that mentions are disambiguated solely based on the phrases themselves. Since the AskAPatient dataset lacks a test set, we employed a 10-fold cross-validation approach as outlined in the original paper by Limsopatham and Collier (2016a). The statistics reported are the averages across these folds.
MM-ST21pv Mohan and Li (2019)
@@
The MedMentions dataset is a large-scale resource for biomedical entity recognition. The ST21pv subset includes 4,392 PubMed abstracts with over 200,000 entity mentions linked to 21 selected UMLS semantic types. This dataset provides a comprehensive resource for training and evaluating biomedical entity recognition systems. Unlike the original dataset, we use the 2020AA version of UMLS as the KBs because the 2017AA version of UMLS is not directly accessible. This leads to some differences after preprocessing due to variations between versions. Specifically, our dataset deviates from the original MedMentions dataset by 741 training samples (0.6%), 284 validation samples (0.7%), and 235 test samples (0.6%).
Appendix C Top-5 Accuracy

### v2

@@ 5.1 Ablation Study
We conducted in-depth analyses on the selection of positive and negative pairs, the effect of similarity models on synonym retrieval, and the effect of pre-training. Additional analyses and results—including the number of synonyms used in positive-only training and the effect of optimization functions—can be found in Appendix C.
Selection of positive and negative pairs
Model | NCBI | BC5CDR |
TF-IDF (trigram-based) | 91.0 | 92.6 |
BioBERT-NLI | 90.1 | 71.9 |
SapBERT | 90.2 | 84.1 |
Analyzing the impact of how positive-negative pairs are constructed during negative-aware training is crucial for determining the optimal strategy for selecting hard negatives and the appropriate number of pairs. We investigated the effects of three factors: (1) negative sampling techniques (i.e., whether to use the model’s incorrect predictions as negatives or rely on TF-IDF-based sampling), (2) the relative ranking of positive and negative samples, and (3) top-k selection (i.e., the number of negatives to include). Detailed results can be found in Table 2. Ultimately, selecting negatives from the model’s incorrect predictions proved to be the most important factor, with an average score difference of 0.8%, while the other factors showed smaller differences of 0.5%. More importantly, regardless of the specific negative-aware training approach, its application leads to significant performance improvements compared to models like GenBioEL, which do not incorporate such training. All models applying negative-aware training, including our ANGEL model, outperformed GenBioEL by 0.9% to 1.7% ( for all comparisons).
Effect of similarity models
Similarity models play a critical role in retrieving synonymous terms from the KB, and their choice can have a substantial impact on overall system performance. To assess their effectiveness, we evaluated three models: (1) TF-IDF, (2) BioBERT-NLI,444https://huggingface.co/gsarti/biobert-nli a sentence embedding model fine-tuned on natural language inference datasets, and (3) SapBERT. These models were incorporated into the positive-aware training framework of GenBioEL on the NCBI-Disease and BC5CDR datasets. As presented in Table 3, the TF-IDF-based approach outperforms the two embedding-based models. Although the strong performance of trigram-based similarity highlights the utility of surface-level matching in BioEL, this does not imply that the task is inherently simple. While many synonyms exhibit similar surface forms, a substantial portion do not—posing challenging edge cases that demand more nuanced semantic understanding.
Effect of pre-training
Model | FT | BC5CDR | AAP |
BART | ✗ | 0.8 | 15.6 |
GenBioEL | ✗ | 33.1 | 50.6 |
+ ANGEL (Ours) | ✗ | 49.7 | 61.5 |
BART | ✓ | 93.0 | 88.7 |
GenBioEL | ✓ | 93.1 | 89.3 |
+ ANGEL (Ours) | ✓ | 94.5 | 90.2 |
@@
Table 4 highlights the effectiveness of ANGEL’s pre-training by comparing other pre-training methods. BART, pre-trained using a standard language modeling objective but not specifically tailored for BioEL tasks, shows the lowest performance. In contrast, GenBioEL, pre-trained using synonyms from UMLS in a similar manner to our positive-only training, initially demonstrates a substantial performance advantage over BART. However, this gap narrows considerably after fine-tuning, to the point where it is no longer statistically significant. When ANGEL’s negative-aware training is applied to GenBioEL, its performance improves significantly, achieving gains of 16.6% on BC5CDR and 10.9% on AAP. Even after fine-tuning, the performance gap remains noticeable, with a difference of 1.4% on BC5CDR and 0.9% on AAP.
5.2 Understanding the Effectiveness of Negative-aware Training
Figure 3 provide an interpretation of how negative-aware training leads to performance improvements. While positive-only training increases the probability of identifying synonyms, it also raises the risk of incorrectly boosting the probability of negative samples that are morphologically similar to the input mention. In contrast, negative-aware training improves the identification of synonyms while simultaneously reducing the probability of incorrect negatives, making it particularly effective when these negatives share morphological similarity with the input mention. To verify this, we divided input mention-negative pairs from the NCBI-disease dataset into 10 bins based on their tri-gram TF-IDF similarity. We then computed the log probabilities of negatives within each bin for the corresponding input mentions, comparing them to the log probabilities of the top-1 positive predictions (i.e., the synonym assigned the highest probability by the model). As the similarity between the input mention and the negative entities increased, the probabilities assigned by GenBioEL to positive and negative samples became more similar, eventually leading to higher probabilities for the negative samples. In contrast, our model demonstrated a clear distinction in behavior, consistently prioritizing positive samples over negatives.
5.3 Error Analysis
We conducted an in-depth evaluation of the models based on the similarity between the input mentions and the gold-standard entities. Similarity was calculated using tri-gram TF-IDF, with the gold-standard entity determined as the candidate synonym with the highest similarity score to the input mention. The similarity scores, ranging from 0 to 1.0, were divided into five bins, and accuracy was measured for each bin. As shown in Figure 4, errors predominantly occurred in the 0-0.2 and 0.2–0.4 bins, as indicated by the height of the hatched bars, which represent the number of errors. This suggests that models tend to struggle when the surface forms of the input mentions are not closely aligned with those of the gold-standard entities. Our method improves the generalizability of the model, leading to an overall reduction in GenBioEL’s errors across all bins, with particularly notable improvements in cases of low similarity. However, significant challenges remain, as the accuracy of our model is only 34.2% in the 0–0.2 bin, highlighting the need for further improvement.
@@ 5.4 Case Study
Figure 5 illustrates the predictions of SapBERT, GenBioEL, and ANGEL. In the first example, the mention ‘ASPD’ is an abbreviation for ‘antisocial personality disorder’ (also known as ‘dissocial personality disorder’). SapBERT incorrectly predicts ‘ASP’ due to the similarity in surface form. GenBioEL struggles to distinguish between correct entity names and those containing the words ‘personality disorder’. In contrast, our model successfully identifies the correct entities, without being misled by false entity names that contain overlapping terms. The second example involves the mention ‘basaglar,’ a medication that contains insulin glargine, a long-acting insulin. The challenge here arises from the fact that product names can differ significantly from the biomedical terms used to describe their active ingredients. This discrepancy leads to failures in both SapBERT and GenBioEL, as they struggle to connect the brand name to its corresponding biomedical entity. Nevertheless, our model successfully identifies the correct entity, showcasing its ability to handle such complex cases effectively. In the final example, our method was less effective. For the mention of ‘glutamine,’ neither SapBERT nor GenBioEL identified the correct answer, but they did rank ‘Glutamic acid,’ the correct entity, within the top 5 candidates. Our model, however, ranked the correct answer slightly lower. Consequently, while our model shows a notable improvement in top-1 accuracy, the increase in top-5 accuracy is relatively modest in some datasets. The effectiveness of our method also varies across different datasets. We discuss this limitation in more detail in Appendix E, noting that such cases are an area for further exploration.
@@ 6 Conclusions
In this study, we discussed the importance of negative samples in training generative BioEL models and introduced ANGEL, the first framework in this field to effectively incorporate negative-aware training into a generative model. Our models demonstrated the ability to learn subtle distinctions between entities with similar surface forms and contexts. Experimental results showed that ANGEL outperformed existing similarity-based and generative models, with notable performance improvements of 0.9%, 1.4%, and 1.7% for BART, BioBART, and GenBioEL, respectively, achieving the best performance across five public BioEL datasets.
Limitations
Our method is versatile and applicable to any generative model, but it has only been tested on encoder-decoder models and not on decoder-only models such as BioGPT Luo et al. (2022). We plan to further investigate the effect of our method on these models. Additionally, it has not been tested on recent open-source large language models (LLMs) Touvron et al. (2023); Chen et al. (2023). While we acknowledge that incorporating comparisons with LLMs and further assessing the effectiveness of our approach would be an interesting direction, using LLMs for entity linking presents new challenges. The primary concern with larger models is their inefficiency, particularly regarding slower inference speeds and higher memory requirements, which may render them unsuitable for most real-world applications. This issue becomes particularly problematic in biomedical information extraction, where processing millions of publications to extract meaningful insights is essential.
Our negative-aware training method may not be limited to a specific domain, yet we have only evaluated it on biomedical-domain datasets, which restricts the demonstration of its broad applicability. Nevertheless, we would like to emphasize the reasons for focusing on the biomedical domain. Biomedical entity linking has unique characteristics that differentiate it from other domains, making this problem both challenging and interesting. In general domains, ambiguity typically arises between different types of entities (e.g., whether “Liverpool” refers to a city or a sports club). Similarly, in the biomedical domain, ambiguity exists between different types, such as whether “Ebola” in Figure 1 refers to a disease or a virus. Additionally, biomedical entities often exhibit significant variations in their surface forms, even when they share the same identifier, i.e., they refer to the same entity. As shown in Figure 5, “Basaglar” can be expressed as other variations such as “insulin glargine substance” or “insulin glargine.” Furthermore, terms like “substance” in the entity “insulin glargine substance” overlap with many other entities (e.g., “Basiliximab substance,” “Linagliptin substance,” “Benzodiazepine substance”), making the task even more complex. Therefore, distinguishing between numerous candidates with similar surface forms is especially crucial in biomedical entity linking. We believe that our method, which trains the model using negative samples with similar structures, is particularly well-suited to tackle this challenge. However, exploring the application of our approach in other domains would be a valuable direction for future research.
Ethical Considerations
This study complies with ethical standards, ensuring that all datasets and models adhere to their respective licenses and usage terms. While our method was evaluated on five widely used datasets, these serve primarily as benchmarks and may not fully capture real-world complexities. Although the model demonstrates significant improvements, its limitations in handling low-similarity cases highlight the need for thorough validation before deployment, particularly in sensitive applications.
@@ Acknowledgements
This research was supported by (1) the National Research Foundation of Korea (NRF-2023R1A2C3004176, RS-2023-00262002), (2) the Ministry of Health & Welfare, Republic of Korea (HR20C002103), (3) ICT Creative Consilience Program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-2025-RS-2020-II201819). M.S. was supported by (1) No. RS-2022-00155911: Artificial Intelligence Convergence Innovation Human Resources Development(Kyung Hee University), (2) No. RS-2024-00509257: Global AI Frontier Lab), and (3) the MSIT(Ministry of Science and ICT), Korea, under the ITRC(Information Technology Research Center) support program(IITP-2024-RS-2024- 00438239, 15%).
Appendix A Datasets
@@
Entity names | 108,092 | 809,929 | 904,798 | 3,398 | 6,051,091 |
Identifiers | 14,944 | 268,162 | 350,830 | 1,036 | 3,092,324 |
@@
Table A presents the statistics of the five datasets used, along with their corresponding target knowledge bases.
NCBI-disease
@@
The NCBI-disease dataset Doğan et al. (2014) contains 793 PubMed abstracts annotated with 6,892 disease mentions that are mapped to 790 unique disease concepts using the MEDIC ontology Davis et al. (2012). MEDIC is a medical dictionary that integrates disease concepts, synonyms, and definitions from both MeSH Lipscomb (2000) and OMIM Hamosh et al. (2004), encompassing a total of 9,700 unique disease entities. This dataset is primarily used for disease recognition and concept normalization tasks.
BC5CDR
@@
The BC5CDR dataset Li et al. (2016b) includes 1,500 PubMed abstracts with 4,409 chemical entities, 5,818 disease entities, and 3,116 chemical-disease interactions. All annotated entities are mapped to the MeSH ontology Lipscomb (2000), which is a subset of UMLS Bodenreider (2004). This dataset is widely used for biomedical entity recognition and interaction studies. To fit the purpose of our study, we use only the chemical and disease annotations and discard the interaction annotations.
COMETA
@@
COMETA Basaldella et al. (2020) focuses on layman medical terminology, compiled from four years of content across 68 health-related subreddits. This dataset consists of 20K biomedical entity mentions annotated with concepts from SNOMED CT Chang and Mostafa (2021). It is utilized for the normalization of consumer health expressions into standardized terminologies.
AskAPatient (AAP)
@@
The AskAPatient dataset Limsopatham and Collier (2016b) contains 8,662 phrases from social media language, each mapped to medical concepts from SNOMED CT Chang and Mostafa (2021). This dataset does not include contextual information, meaning that mentions are disambiguated solely based on the phrases themselves. Since the AskAPatient dataset lacks a test set, we employed a 10-fold cross-validation approach as outlined in the original paper by Limsopatham and Collier (2016a). The statistics reported are the averages across these folds.
MM-ST21pv
@@
The Medmentions dataset Mohan and Li (2019) is a large-scale resource for biomedical entity recognition. The ST21pv subset includes 4,392 PubMed abstracts with over 200,000 entity mentions linked to 21 selected UMLS semantic types. This dataset provides a comprehensive resource for training and evaluating biomedical entity recognition systems. Unlike the original dataset, we use the 2020AA version of UMLS as the KBs because the 2017AA version of UMLS is not directly accessible. This leads to some differences after preprocessing due to variations between versions. Specifically, our dataset deviates from the original Medmentions dataset by 741 training samples (0.6%), 284 validation samples (0.7%), and 235 test samples (0.6%).
Appendix B Hyperparameter Configurations
Hyperparameter | Pre-training | Fine-tuning | ||||
NCBI | BC5CDR | COMETA | AAP | MM-ST21pv | ||
Positive-only Training | ||||||
Training Steps | 80K | 20K | 30K | 40K | 30K | 40K |
Learning Rate | 4e-5 | 3e-7 | 5e-6 | 2e-5 | 5e-6 | 3e-5 |
Weight Decay | 0.01 | 0.01 | 0.01 | 0.01 | 0.01 | 0.01 |
Batch Size | 384 | 16 | 16 | 16 | 16 | 16 |
Adam | 1e-8 | 1e-8 | 1e-8 | 1e-8 | 1e-8 | 1e-8 |
Warmup Steps | 1,600 | 0 | 500 | 1000 | 0 | 1,000 |
Attention Dropout | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 |
Clipping Grad | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 |
Label Smoothing | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 |
Negative-aware Training | ||||||
Epochs | 5 | 1 | 1 | 1 | 1 | 1 |
Learning Rate | 1e-5 | 1e-5 | 1e-6 | 5e-6 | 5e-6 | 5e-6 |
(DPO) | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 |
Weight Decay | 0.01 | 0.01 | 0.01 | 0.01 | 0.01 | 0.01 |
Batch Size | 64 | 16 | 16 | 32 | 8 | 16 |
Warmup Steps | 1000 | - | - | - | - | - |
Table B details the hyperparameters used for positive-only training and negative-aware training across the BioEL benchmark datasets. We searched for the optimal hyperparameter settings using the validation sets. We refer to the study of Yuan et al. (2022b) to determine the range of the hyperparameters. During pre-training, we used the same hyperparameters as in GenBioEL. For positive-only training, we explored a range of training steps between 20K and 40K, a learning rate between 2e-5 and 3e-7, and batch sizes from 8 to 16, except during pre-training. During negative-aware training, we fixed the at 0.1, in accordance with the basic configuration of DPO, and searched the hyperparameter space using a learning rate between 2e-5 and 1e-6 and batch sizes ranging from 8 to 64. We used the source codes provided by Yuan et al. (2022b)555https://github.com/Yuanhy1997/GenBioEL and alignment handbook Tunstall et al. (2023)666https://github.com/huggingface/alignment-handbook.
Appendix C Ablation Study
Model | FT | NCBI | COMETA | MM |
BART | ✗ | 10.7 | 8.4 | 0.9 |
GenBioEL | ✗ | 58.2 | 42.4 | 10.4 |
+ ANGEL (Ours) | ✗ | 64.6 | 49.8 | 18.2 |
BART | ✓ | 90.3 | 80.4 | 70.1 |
GenBioEL | ✓ | 91.0 | 80.9 | 70.7 |
+ ANGEL (Ours) | ✓ | 92.8 | 82.8 | 73.3 |
Effect of pre-training
Effect of optimization functions
Model | Acc@1 |
GenBioEL | 85.0 |
ANGELFT (Pairwise) | 85.9 |
ANGELFT (CPO) Xu et al. (2024) | 85.9 |
ANGELFT (SimPO) Meng et al. (2024) | 86.1 |
ANGELFT (DPO) Rafailov et al. (2024) | 86.2 |
Our negative-aware framework is compatible with various optimization methods. To demonstrate this flexibility, we evaluated three additional loss functions during fine-tuning: (i) a simple pairwise loss, where is defined as ; and two preference optimization methods that build upon and improve DPO: (ii) Contrastive Preference Optimization (CPO) Xu et al. (2024), and (iii) Similarity Preference Optimization (SimPO) Meng et al. (2024). Note that Equation 4 in the Method section presents a simplified version for clarity of explanation. In practice, CPO and SimPO introduce additional terms and require slight extensions to this formulation. For a detailed comparison of the exact equations, we refer readers to Table 7 in Meng et al. (2024), which provides a clear summary of the differences.
As shown in Table D, while DPO achieves the highest accuracy, the performance differences among the optimizers are relatively small. Regardless of the specific method used, all optimizers consistently enhance performance and outperform the GenBioEL baseline.
The number of synonyms
To evaluate the impact of incorporating multiple synonyms during fine-tuning (Equation 3), we conducted experiments by varying the number of synonyms associated with each mention, testing with 1, 3, and 5 synonyms. As a result, using 3 synonyms proved to be optimal, outperforming the approach that used only a single top-1 synonym in the study of Yuan et al. (2022b).
Appendix D Error Analysis
Consistent with the analysis on the NCBI-disease dataset (Figure 4), Figure B reveals that models on the COMETA dataset most frequently made errors in the 0.0–0.2 bin, where input mentions have low similarity to gold-standard entities. Across all similarity bins, our ANGEL framework consistently improved upon GenBioEL’s performance, leading to overall gains. A similar trend is observed in Figure C for the MedMentions dataset, where ANGEL again outperforms GenBioEL across all bins. These results highlight the need for future research focused on reducing errors in low-similarity scenarios.
Appendix E Top-5 Accuracy